OpenStudy (mrnebula):

Which best describes your ability to solve quadratic equations? *I can use multiple strategies to solve quadratic equations and I can teach someone else how to solve them as well. *I can use multiple strategies to solve quadratic equations, but sometimes I make mistakes. *I can use at least one strategy to solve quadratic equations, but I make some mistakes. *I do not understand how to solve quadratic equations.

OpenStudy (mrnebula):

@undeadknight26 @Hotchellerae21 @welshfella

Nnesha (nnesha):

you should answer that question they are asking how much u know abt quadratic

OpenStudy (mrnebula):

But what if u dont know i just ot back into algebra i dont remember anything

OpenStudy (mrnebula):

I guess i just guess

Nnesha (nnesha):

then you should pick last one

OpenStudy (mrnebula):

Oh ok thanks

Nnesha (nnesha):

\[\huge\rm Ax^2+Bx+C=0\] quadratic equation where a-leading coefficient b-middle term c-constant term formula \[\huge\rm \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] there are 3+ ways to solve qudratic equation factoring graphing qudratic formula AC method

Nnesha (nnesha):

you can learn how to solve quadratic equation and then u can pick 3